[Hallc_running] [Revised Logentry] RC Daily Update

Xiaochao Zheng xiaochao at jlab.org
Thu Aug 27 13:24:00 EDT 2020

Hi Burcu:

Thanks for the report. I just want to point out two things that I noticed when reading the report:

By 1:30am, the reference cell was filled with PolHe3 (https://logbooks.jlab.org/entry/3835347).

While we were getting ready for an access and during the escorted access, TO evacuated the reference cell

1) the ref cell has to be fileld with regular He3 gas, not PolHe3.
2) Do I read correctly that we purged the 3He gas only an hour after filling the ref cell?  Was it an accident or intentional?



From: Hallc_running <hallc_running-bounces at jlab.org> on behalf of bduran at jlab.org <bduran at jlab.org>
Sent: Thursday, August 27, 2020 12:20 PM
To: hallc_running at jlab.org <hallc_running at jlab.org>
Subject: [Hallc_running] [Revised Logentry] RC Daily Update

RC Daily Update<https://logbooks.jlab.org/entry/3835654>

  *   Edit<https://logbooks.jlab.org/entry/3835654/edit?destination=email/send>
  *   Delete<https://logbooks.jlab.org/entry/3835654/delete?destination=email/send>

Lognumber 3835654<https://logbooks.jlab.org/entry/3835654>. Submitted by bduran<https://logbooks.jlab.org/user/bduran> on Thu, 08/27/2020 - 11:52<https://logbooks.jlab.org/entries?start_date=1598539945&end_date=1598547145&book=HCLOG>.

Last updated on Thu, 08/27/2020 - 12:18

Logbooks:       HCLOG<https://logbooks.jlab.org/book/hclog>
Entry Makers:   bduran

Yesterday at the beginning of the swing shift, MCC had to work a bit on tuning the beam for us and did harp scans. Following that we resumed production running with PolHe3 longitudinal and we took data until 6pm when Brad came in to test new (correct) beam positions for glass cells (https://logbooks.jlab.org/entry/3835151). While he was testing the new beam positions, we had an interlock on HMS Q2 on LHe level (~30%). Heidi stopped by the CH and informed us the problem is corrected and asked us to wait the LHe levels reach 35-40% to ramp the magnet up. At 8:04pm, the magnet was up again (https://logbooks.jlab.org/entry/3835190). During the new beam positions testing (we use 5 uA CW or tune), MCC was struggling to send us beam and they reported the Compton ion chamber was acting up in a way similar to the transport ion chamber (they're reading higher signals with no beam or lower currents).

We wrapped up the production running with PolHe3 longitudinal by owl shift yesterday. The last NMR measurement @8:28 pm : 44.1% ( https://logbooks.jlab.org/entry/3835209) and PNMR: 44.3% (https://logbooks.jlab.org/entry/3835225)

At 12:30am we took an escorted access to purge and fill the reference cell with He3. This was needed to carry out taking data with reference cell He3 while the target experts prepare the PolHe3 target for transverse running. By 1:30am, the reference cell was filled with PolHe3 (https://logbooks.jlab.org/entry/3835347).

While we were getting ready for an access and during the escorted access, TO evacuated the reference cell and started the AFP loss and temperature tests that were needed to be complete before rotating the PolHe3 to transverse. At 2:30am, the tests were done (https://logbooks.jlab.org/entry/3835396), target lasers are switched to transverse direction (https://logbooks.jlab.org/entry/3835377) and the target spin direction was rotated to 270 degrees (https://logbooks.jlab.org/entry/3835381).
Around 2:30am we had about half an hour down time due to hall C dipole magnet MBSY3C string tripped off (https://logbooks.jlab.org/entry/3835410). Following that, we took production data with reference cell (He3) until 06:43am when Mingyu/Junhao began the EPR/NMR calibration procedure. At 08:24am we resumed data taking with the reference cell (He3).

Since 10:55am, we have been down for an escorted access to complete the following tasks:

1) Troubleshooting the transport ion chamber issue (https://logbooks.jlab.org/entry/3835254), Jerry Kowal's team is taking care of this. (ongoing)

2) Reset rotation drive controllers (Joe Beaufait) (done)

3) Connecting the spare air compressor that Walter rented. (https://logbooks.jlab.org/entry/3833255) (Larry et al) (done)

4) Tweaking the EPR RF amplifier gain setting to troubleshoot the noisy NMR (https://logbooks.jlab.org/entry/3835639) (done)

5) Check air flow/fan in laser box ? (Brad)

We expect to be down about ~ 2hours.

We will have an RC Meeting at 4:15 pm

Bluejeans remote connection information:
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial: +1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200827/5931b8f3/attachment-0001.html>

More information about the Hallc_running mailing list